if tanαtanβ =-a^2/b^2 then the chord joining α and β will subtend right angles at(I) Focus(II) Centre(III)Major Axis(IV)Minor axis

Grade:11

1 Answers

Dilli PRASAD
6 years ago
Given tan h . tank =-a^2/b^2b^2/a^2). Tanh .tank =-1(bsinh/acosh)(bsink/acosk)=-1(bsinh-0)/(acosh-0)*(bsink-0)/(acosk-0)=-1CP slope * CQ slope =-1 whare C(0,0) P(acosh,bsinh) Q(acosk,bsink)There fore PQ makes right angle at C